Consider a gas that occupies 1.00 dm3 at a pressure of 2.00 bar. If the gas is compressed isothermally at constant external pressure, Pext, so that the final volume is 0.500 dm3 , what is the smallest value Pext can have? Calculate the work involved using this value of Pext.


1
Expert's answer
2021-09-04T02:28:54-0400

P1V1 =P2V2

1× 2 = 0.5 × P2

P2 = 2 × 2 = 4 bar